A series RL circuit with L = 3.00 H and a series RC circuit with C = 3.00 F have equal time constants. If the two circuits contain the same resistances R, (a) what is the value of R and (b) what sit the time constant?
Answer #1
A series RL circuit with inductance L = 3.00 H series with RC circuit
capacitance C = 3.00 F
   resistance is R
a ) time constant for RC circuit is τ = RC  ........(1)
time constant for RL circuit is τ = L / R ........(2)
from the above two R^2 = L / C
                                 R = √ L/ C
                                       = √ 3.00 H / 3.00 F
                                        = 1 Ω
b ) time constant is τ = RC
                                     = 1.00Ω * 3.00 F
                                     = 3.00 s